# Large Language Models for Cyber Security: A Systematic Literature Review

### v2

@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Botacin, 2023; Liu et al., 2023e). LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email content and identifying deceptive language patterns (Jamal and Wimmer, 2023; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2023). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
While there have been several valuable efforts in the literature to survey the LLM4Security (Yao et al., 2024; de Jesus Coelho da Silva and Westphall, 2024; Motlagh et al., 2024; Divakaran and Peddinti, 2024) given the growing body of work in this direction, these studies often have a more focused scope. Many of the existing surveys primarily concentrate on reviewing the types of tasks that LLMs can be applied to, without providing an extensive analysis of other essential aspects related to these tasks, such as the data and domain-specific techniques employed (Yigit et al., 2024; Novelli et al., 2024), as shown in Table 1. For example, Divakaran et al. (Divakaran and Peddinti, 2024) only analyzed the prospects and challenges of LLMs in various security tasks, discussing the characteristics of each task separately. However, it lacks insight into the connection between the requirements of these security tasks and data, as well as the application of LLMs in domain-specific technologies.
To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 38,112 relevant paperss and systematically analyzing 127 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
  • RQ1: What LLMs have been employed to support security tasks?
  • RQ2: What types of security tasks have been facilitated by LLM-based approaches?
@@
  • RQ4: What is the difference in data collection and pre-processing when applying LLMs to security tasks?
For each research question, we provide a fine-grained analysis of the approaches, datasets, and evaluation protocols used in the surveyed papers. We identify common themes and categorize the papers along different dimensions to provide a structured overview of the landscape. Furthermore, we highlight the key challenges and limitations of current approaches to guide future research towards addressing the gaps. We believe our survey can serve as a valuable resource for researchers working at the intersection of NLP, AI, and cybersecurity. The contributions of this work are summarized as follows:
  • We conduct a comprehensive Systematic Literature Review (SLR) to investigate the latest research on LLM4Security, providing a mapping of the current landscape. Our search covers an extensive number of over 38,112 papers, with further quality-based and relevance-based filtering, we retain 127 papers for later detailed review.
@@
The survey progresses with the following framework. We outline our survey methodology, which includes the search strategy, inclusion/exclusion criteria, and data extraction process, in Section 2. The analysis and findings for each of the four research questions can be found in Sections 3 through 6. Sections 7 to 8 explore the constraints and significance of our results, while also identifying promising directions for future research. Finally, Section 9 concludes the paper.
@@ 2. Methodology
Planning. Initially, we formulated precise research questions to understand how LLMs are being utilized in security tasks, the benefits derived, and the associated challenges. Subsequently, we developed a detailed protocol delineating our search strategy, including specific databases and keywords, inclusion and exclusion criteria, and data extraction. Each co-author reviewed this protocol to enhance its robustness and align with our research objectives.
Conducting. We meticulously crafted our literature search to ensure comprehensiveness, employing both manual and automated strategies across various databases to encompass a wide range of studies. Each study identified underwent a stringent screening process, initially based on title and abstract, followed by a thorough review of the full text to ensure conformity with our predefined criteria. To prevent overlooking related papers, we also conducted forward and backward snowballing on the collected papers.
@@ 2.1. Research Question
  • RQ1: What LLMs have been employed to support security tasks? This RQ seeks to inventory the specific LLMs that have been utilized in security tasks. Understanding the variety and characteristics of LLMs used can offer insights into their versatility and suitability for different security applications. We will discuss the architectural differences of LLMs and delve into analyzing the impact of LLMs with different architectures on cybersecurity research over different periods.
  • RQ2: What types of security tasks have been facilitated by LLM-based approaches? Here, the focus is on the scope and nature of security tasks that LLMs have been applied to. The goal is to categorize and understand the breadth of security challenges that LLMs are being used to address, highlighting the model’s adaptability and effectiveness across various security dimensions. We will categorize previous studies according to different security domains and provide detailed insights into the diverse security tasks applied by LLMs in each security domain.
  • RQ3: What domain specification techniques are used to adapt LLMs to security tasks? This RQ delves into the specific methodologies and techniques employed to fine-tune or adapt LLMs for security tasks. Understanding these techniques can provide valuable insights into the customization processes that enhance LLM effectiveness in specialized tasks. We will elucidate how LLMs are applied to security tasks by analyzing the domain-specific techniques employed in papers, uncovering the inherent and specific connections between these techniques and particular security tasks.
@@ 2.2. Search Strategy
Step1: Identify related venues and databases. To initiate this approach, we first identify specific venues for manual search and then choose suitable engines (libraries and databases) for automated search. In this stage, we opt for six of the top Security conferences and journals (i.e., S&P, NDSS, USENIX Security, CCS, TDSC, and TIFS) as well as six of the leading Software Engineering conferences and journals (i.e.,ICSE, ESEC/FSE, ISSTA, ASE, TOSEM, and TSE). Given the emerging nature of LLMs in research, we also include arXiv in both manual and automated searches, enabling us to capture the latest unpublished studies in this rapidly evolving field. For automated searches, we select seven widely utilized databases, namely the ACM Digital Library, IEEE Xplore, Science Direct, Web of Science, Springer, Wiley, and arXiv. These databases offer comprehensive coverage of computer science literature and are commonly employed in systematic reviews within this domain (Hou et al., 2024; Zhan et al., 2021; Zhou et al., 2024).
Step2: Establish QGS. Following the completion of screening and reaching agreement on the selection criteria, we manually search the chosen venues to identify all pertinent papers. A total of 41 papers are manually identified during this step, aligning with the research objective and encompassing various techniques, application domains, and evaluation methods.
Step3: Define search keywords. The keywords for automatic search are elicited from articles in the QGS through word frequency or content analysis tools. The search string consists of two sets of keywords:
@@
Step4: Conduct automated search. These identified keywords are connected by logic operators and input to automated search across the above-mentioned seven widely used databases. Our automated search focused on papers published after 2019, in which GPT-2 was published, as it marked a significant milestone in the development of large language models. The search was conducted in the title, abstract, and keyword fields of the papers in each database. Specifically, the number of papers retrieved from each database after applying the search query and the year filter (2019-2023) is as follows: 3,398 papers in ACM Digital Library, 2,112 papers in IEEE Xplore, 724 papers in Science Direct, 4,245 papers in Web of Science, 23,721 papers in Springer, 7,154 papers in Wiley, and 3,557 papers in arXiv.
@@ 2.3. Study Selection
In#1: The title and abstract of the paper clearly contain the search keywords;
In#2: Full research papers published in identified conferences or journals, or as preprints on arXiv;
In#3: Papers that apply large language models (e.g., BERT, GPT, T5) to security tasks;
In#4: Papers that propose new techniques or models for security tasks based on large language models;
In#5: Papers that evaluate the performance or effectiveness of large language models in security contexts.
@@
Ex#3: Papers that do not focus on security tasks (e.g., natural language processing tasks in general domains);
Ex#4: Papers that use traditional machine learning or deep learning techniques without involving large language models;
Ex#5: Secondary studies, such as an SLR, review, or survey;
Ex#6: Papers not written in English;
Ex#7: Papers focus on the security of LLMs rather than using LLM for security tasks.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined five inclusion criteria and seven exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, In#2, Ex#1, and Ex#2 were automatically filtered based on the keywords, publication venue, length, and duplication status of the papers. The remaining inclusion criteria (In#3~5) and exclusion criteria (Ex#3~7) were manually applied by inspecting the topic and content of each paper. Specifically, In#1 retained 7,582 papers whose titles and abstracts clearly contained the search keywords. Subsequently, Ex#1 filtered out 440 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 4,855 papers being removed. The remaining papers were then manually screened based on the criteria In#2, which retained 523 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#3~5, Ex#3~7) were then manually applied to the titles and abstracts of these 523 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 156 papers were included for further fine-grained full-text quality assessment.
@@ 2.3.3. Forward and Backward Snowballing
We applied the same inclusion/exclusion criteria and quality assessment to the papers found through snowballing. After the iterations of snowballing, 44 additional papers were included, resulting in a final set of 127 papers for data extraction and synthesis.
@@ 2.4. Statistics of Selected Papers
After conducting searches and snowballing, a total of 127 relevant research papers were ultimately obtained. The distribution of the included documents is outlined in Figure 3. As depicted in Figure 3(A), 36% of the papers underwent peer review before publication. Among these venues, ICSE had the highest frequency, contributing 7%. Other venues making significant contributions included FSE, ISSTA, ASE, and TSE, with contributions of 5%, 5%, 3%, and 3% respectively. Meanwhile, the remaining 62% of the papers were published on arXiv, an open-access platform serving as a repository for scholarly articles. This discovery is unsurprising given the rapid emergence of new LLM4Security studies, with many works recently completed and potentially undergoing peer review. Despite lacking peer review, we conducted rigorous quality assessments on all collected papers to ensure the integrity of our investigation results. This approach enables us to include all high-quality and relevant publications while upholding stringent research standards.
@@
2,3,4 The category of cybersecurity domain
@@
2,3 Security task requirements and the application of LLM solutions

### v3

@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Joyce et al., 2023; Liu et al., 2023e). LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email contents and identifying deceptive language patterns (Jamal and Wimmer, 2023; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2023). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
While there have been several valuable efforts in the literature to survey the LLM4Security (Yao et al., 2024; de Jesus Coelho da Silva and Westphall, 2024; Motlagh et al., 2024; Divakaran and Peddinti, 2024), given the growing body of work in this direction, these studies often have a more focused scope. Many of the existing surveys primarily concentrate on reviewing the types of tasks that LLMs can be applied to, without providing an extensive analysis of other essential aspects related to these tasks, such as the data and domain-specific techniques employed (Yigit et al., 2024; Novelli et al., 2024), as shown in Table 1. For example, Divakaran et al. (Divakaran and Peddinti, 2024) only analyzed the prospects and challenges of LLMs in various security tasks, discussing the characteristics of each task separately. However, it lacks insight into the connection between the requirements of these security tasks and data, as well as the application of LLMs in domain-specific technologies.
To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 38,112 relevant papers and systematically analyzing 127 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
  • RQ1: What types of security tasks have been facilitated by LLM-based approaches?
  • RQ2: What LLMs have been employed to support security tasks?
@@
  • RQ4: What is the difference in data collection and pre-processing when applying LLMs to various security tasks?
For each research question, we provide a fine-grained analysis of the approaches, datasets, and evaluation methodologies used in the surveyed papers. We identify common themes and categorize the papers along different dimensions to provide a structured overview of the landscape. Furthermore, we highlight the key challenges and limitations of current approaches to guide future research towards addressing the gaps. We believe our survey can serve as a valuable resource for researchers working at the intersection of NLP, AI, and cybersecurity. The contributions of this work are summarized as follows:
  • We conduct a comprehensive Systematic Literature Review (SLR) to investigate the latest research on LLM4Security, providing a mapping of the current landscape. Our search covers an extensive number of over 38,112 papers. With further quality-based and relevance-based filtering, we retain 127 papers for later detailed review.
@@
The survey progresses with the following framework. We outline our survey methodology, including the search strategy, inclusion/exclusion criteria, and the data extraction process, in Section 2. The analysis and findings for each of the four research questions can be found in Sections 4 through 6. Sections 7 to 8 explore the constraints and significance of our results, while also identifying promising directions for future research. Finally, Section 9 concludes the paper.
@@ 2. Methodology
Planning. Initially, we formulated precise research questions to understand how LLMs are being utilized in security tasks, the benefits derived, and the associated challenges. Subsequently, we developed a detailed protocol delineating our search strategy, including specific venues and databases, keywords, and quality assessment criteria. Each co-author reviewed this protocol to enhance its robustness and align with our research objectives.
Literature survey and analysis. We meticulously crafted our literature search to ensure comprehensiveness, employing both manual and automated strategies across various databases to encompass a wide range of papers. Each study identified underwent a stringent screening process, initially based on their titles and abstracts, followed by a thorough review of the full text to ensure conformity with our predefined criteria. To prevent overlooking related papers, we also conducted forward and backward snowballing on the collected papers.
@@ 2.1. Research Question
  • RQ1: What types of security tasks have been facilitated by LLM-based approaches? Here, the focus is on the scope and nature of security tasks that LLMs have been applied to. The goal is to categorize and understand the breadth of security challenges that LLMs are being used to address, highlighting the model’s adaptability and effectiveness across various security dimensions. We will categorize previous studies according to different security domains and provide detailed insights into the diverse security tasks that use LLMs in each security domain.
  • RQ2: What LLMs have been employed to support security tasks? This RQ seeks to inventory the specific LLMs that have been utilized in security tasks. Understanding the variety and characteristics of LLMs used can offer insights into their versatility and suitability for different security applications. We will discuss the architectural differences of LLMs and delve into analyzing the impact of LLMs with different architectures on cybersecurity research over different periods.
  • RQ3: What domain specification techniques are used to adapt LLMs to security tasks? This RQ delves into the specific methodologies and techniques employed to fine-tune or adapt LLMs for security tasks. Understanding these techniques can provide valuable insights into the customization processes that enhance LLMs’ effectiveness in specialized tasks. We will elucidate how LLMs are applied to security tasks by analyzing the domain-specific techniques employed in papers, uncovering the inherent and specific connections between these techniques and particular security tasks.
@@ 2.2. Search Strategy
Step1: Identify related venues and databases. To initiate this approach, we first identify specific venues for manual search and then choose suitable libraries and databases for the automated search. In this stage, we opt for six of the top Security conferences and journals (i.e., S&P, NDSS, USENIX Security, CCS, TDSC, and TIFS) as well as six of the leading Software Engineering conferences and journals (i.e.,ICSE, ESEC/FSE, ISSTA, ASE, TOSEM, and TSE). Given the emerging nature of LLMs in research, we also include arXiv in both manual and automated searches, enabling us to capture the latest unpublished studies in this rapidly evolving field. For automated searches, we select seven widely utilized databases, namely the ACM Digital Library, IEEE Xplore, Science Direct, Web of Science, Springer, Wiley, and arXiv. These databases offer comprehensive coverage of computer science literature and are commonly employed in systematic reviews within this domain (Hou et al., 2024; Zhan et al., 2021; Zhou et al., 2024).
Step2: Establish QGS. In this step, we start with creating a manually curated set of studies that have been carefully screened to form the QGS. A total of 41 papers relevant to LLM4Sec are manually identified, aligning with the research objective and encompassing various techniques, application domains, and evaluation methods.
Step3: Define search keywords. The keywords for automatic search are elicited from the title and abstract of the selected QGS papers through word frequency analysis. The search string consists of two sets of keywords:
@@
Step4: Conduct an automated search. These identified keywords are paired one by one and input into automated searches across the above-mentioned seven widely used databases. Our automated search focused on papers published after 2019, in which GPT-2 was published, as it marked a significant milestone in the development of large language models. The search was conducted in the title, abstract, and keyword fields of the papers in each database. Specifically, the number of papers retrieved from each database after applying the search query and the year filter (2019-2023) is as follows: 3,398 papers in ACM Digital Library, 2,112 papers in IEEE Xplore, 724 papers in Science Direct, 4,245 papers in Web of Science, 23,721 papers in Springer, 7,154 papers in Wiley, and 3,557 papers in arXiv.
@@ 2.3. Study Selection
In#1: The title and abstract of the paper contain a pair of identified search keywords;
In#2: Papers that apply large language models (e.g., BERT, GPT, T5) to security tasks;
In#3: Papers that propose new techniques or models for security tasks based on large language models;
In#4: Papers that evaluate the performance or effectiveness of large language models in security contexts.
@@
Ex#3: Papers not published in identified conferences or journals, nor as preprints on arXiv; Ex#4: Papers that do not focus on security tasks (e.g., natural language processing tasks in general domains);
Ex#5: Papers that use traditional machine learning or deep learning techniques without involving large language models;
Ex#6: Secondary studies, such as an SLR, review, or survey;
Ex#7: Papers not written in English;
Ex#8: Papers focus on the security of LLMs rather than using LLMs for security tasks.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined four inclusion criteria and eight exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, Ex#1, Ex#2, and Ex#3 were automatically filtered based on the keywords, duplication status, length, and publication venue of the papers. The remaining inclusion criteria (In#2~4) and exclusion criteria (Ex#4~8) were manually applied by inspecting the topic and content of each paper. Specifically, the criteria of In#1 retained 7,582 papers whose titles and abstracts contained a pair of the identified search keywords. Subsequently, Ex#1 filtered out 440 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 4,855 papers being removed. The remaining papers were then screened based on the criteria Ex#3, which retained 523 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#2~4, Ex#4~8) were then manually applied to the titles and abstracts of these 523 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 156 papers were included for further fine-grained full-text quality assessment.
@@ 2.3.3. Forward and Backward Snowballing
Here we obtained 2,056 and 5,255 papers separately during the forward and backward process. Then we applied the same inclusion/exclusion criteria and quality assessment to the papers found through snowballing. After the initial keyword filtering and deduplication, there were 1,978 papers that remained available. Among them, 68 papers were excluded during the page number filtering step, and 1,235 papers were deleted to ensure the papers were published in the selected venues. After confirming the paper topics and assessing the paper quality, only 44 papers were ultimately retained in the snowballing process, resulting in a final set of 127 papers for data extraction and synthesis.
@@ 2.4. Statistics of Selected Papers
After conducting searches and snowballing, a total of 127 relevant research papers were ultimately obtained. The distribution of the included documents is outlined in Figure 3. As depicted in Figure 3(A), 39% of the papers underwent peer review before publication. Among these venues, ICSE had the highest frequency, contributing 7%. Other venues making significant contributions included FSE, ISSTA, ASE, and TSE, with contributions of 5%, 5%, 3%, and 3% respectively. Meanwhile, the remaining 61% of the papers were published on arXiv, an open-access platform serving as a repository for scholarly articles. This discovery is unsurprising given the rapid emergence of new LLM4Security studies, with many works recently completed and potentially undergoing peer review. Despite lacking peer review, we conducted rigorous quality assessments on all collected papers to ensure the integrity of our investigation results. This approach enables us to include all high-quality and relevant publications while upholding stringent research standards.
Refer to caption Refer to caption
@@
1,3,4 The category of cybersecurity domain
@@
1,3 Security task requirements and the application of LLM solutions